Systems Staff Meeting Notes 2/7/00

Systems Staff Meeting Notes
February 7, 2000

Present: Rich Bennett, Will Chaney, Bill Covey, Debra Harris, Bobbie Meng, Everett Phillips, Hilton Soto, Ying Tang
  1. Bill's Dilbert calendar is back on his desk.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. No major machine orders have gone out.
  2. A few odd software orders have been made.
  1. No decision yet on CSS (cross site scripting). CSS will embed something on a link that activates the destruction when the link is hit. The only solution we can come up with will cripple the functionality of our network. A list of trusted sites could be developed, but there could be problems keeping it updated. There is no way of deciphering the point of origination of the attack. There are even more problems with the SQL server and Oracle, where a combination of commands can result in all kinds of nasty things. We can try to fix it with multiple security scans, but that is only the equivalent of building walls and fences.
  2. DDOS - Distributed Denial of Service attacks - designed to keep you from using your own network by pouring information into the network, overloading it to where it will not work. Tribal network flood - distributes little "Manchurian candidates" all over the network that are designed to go off at a given time, flooding the main distribution stream. This is difficult, if not impossible, to track. Going back to private networks may be the only solution - and that is not feasible.
  1. CSS & DDOS were discussed at the last meeting.
  2. Time Standards - D4 freeware will keep all machines in sync with the national time standard. It is located on NT2 in the Read Only file. Just follow the instructions -right click on the icon to install it. It can run hidden.
  3. ATC - Bill demoed it at MMM, Winston at Liaisons. Met with enthusiastic response - right now it will be updated once a month. The download time is running about 1-2 calendar days at the moment, but we're working on that with NERDC. We want to automate it to update itself periodically. It is a stripped-down application that runs fast - any added features will slow it down.
  1. Resource Services - timeout problems - Will & crew have been swapping out switches for the hubs. They swapped out the main feed switch on the second floor, this seems to have resolved the timeout problems. The Catalyst 1900 switch that was swapped out has not been performing up to spec, according to NERDC. We'll trade out all building feeds using it immediately and talk to CISCO about trade-in options.
  2. NTAC - current debate for one protocol - IP is what is most widely used. Other protocols include IPX and AppleTalk. The Libraries mainly use IP, only IPX is used to talk to the CD ROM server. Much of the information supplied by the CD ROMs is becoming available online. The Libraries no longer need AppleTalk. In theory, the one protocol standard makes sense, however, in reality - it needs to be a slow process of conversion.
  3. Lana Jamerson should be onboard and ready to work today - starting with the wiring documentation project. She'll be located in the order room, we'll work her schedule around Lindsay's.
  4. West wiring project should kick off soon. First on the agenda will be getting the main conduits to carry trunks.
Software Projects Status
  1. ATC demo in MMM, Liaisons was well received.
  2. Cataloger's Desktop seems to be working again in Resource Services & Documents. Of course, it is due for an upgrade in March.
  3. Hilton has been working on changing Web projects and SQL server patches - also assisting Winston with the Preservation project.
  4. Ying is working on the OPS cost report - has two formats finished, working on the third.
  5. Changes to the IR database should be completed this week.
  1. Winston is out all this week, Feb. 7-11.
  2. Suzy is out Feb. 7-9.
  3. Winston out Feb. 16-18.
  4. Hilton out Feb. 17.

Back to the Systems Homepage

Last updated February 14, 2000
by Debra Harris